There is a 3/4 chance that my trash will be taken out. There is a 1/2 chance that my room will be swept. There is a 3/8 chance that both my trash will be taken out and the room will be swept. What is the probability that at least either my trash will be taken out or my room will be swept.?
Answer by ewatrrr(24785)


Hi
P(A or B) = P(A) + P(B) - P(A and B)
P(T 0r S) = 3/4 + 1/2 - 3/8
= 6/8 + 4/8 - 3/8
= 7/8